OpenStudy (anonymous):

Cynthia has a cup full of dimes and quarters that she would like to put in her savings account. When she takes the cup to the bank, the banker tells Cynthia that the 386 coins in the cup are worth $58.10. How many quarters does Cynthia have in her cup?

OpenStudy (anonymous):

Solve the system of equations dimes +quarters = 386 x+y=386 .10x+.25y=58.10

OpenStudy (anonymous):

ok im doing it now

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

What about

OpenStudy (anonymous):

do i plug it in

OpenStudy (anonymous):

x+y=386 solve for either variable y=386-x .10x+.25(386-x)=58.10

OpenStudy (anonymous):

Is the information you provided for the problem correct because I'm not getting whole numbers

OpenStudy (anonymous):

i did that but its not coming to me

OpenStudy (anonymous):

when i solved the eq I got x= 222.7 coins and y = 143.3 coins, which isn't possible for a money problem

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

Got it...trial and error 256 dimes and 130 quarters 256+130=386 25.60+ 32.50=58.10